Niles Co. has the following data related to an item of inventory: Inventory, March 1 400 units @ $2.10 Purchase, March 7 PurcNiles Co. has the following data related to an item of inventory: Inventory, March 1 400 units @ $2.10 Purchase, March 7 1400 units @ $2.20 Purchase, March 16 280 units @ $2.50 Inventory, March 31 520 units The value assigned to ending inventory if Niles uses LIFO is

*PLEASE EXPLAIN IN DETAIL, HOW DOES EVERY NUMBER COME FROM*
Answer #1

Under Lifo, last purchases are sold first. Hence ending inventory belongs to first available units.

So start from beginning inventory and count to 520 units

= (400*2.10) + (520 - 400)*2.20

= (400*2.10) + (120*2.20)

= 840 + 264

= 1104